# Super-LIO: A Robust and Efficient LiDAR-Inertial Odometry System with a Compact Mapping Strategy

### v1

@@ Abstract
LiDAR-Inertial Odometry (LIO) is a foundational technique for autonomous systems, yet its deployment on resource-constrained platforms remains challenging due to computational and memory limitations. We propose Super-LIO, a robust LIO system that demands both high performance and accuracy, ideal for applications such as aerial robots and mobile autonomous systems. At the core of Super-LIO is a compact octo-voxel-based map structure, termed OctVox, that limits each voxel to eight fused subvoxels, enabling strict point density control and incremental denoising during map updates. This design enables a simple yet efficient and accurate map structure, which can be easily integrated into existing LIO frameworks. Additionally, Super-LIO designs a heuristic-guided KNN strategy (HKNN) that accelerates the correspondence search by leveraging spatial locality, further reducing runtime overhead. We evaluated the proposed system using four publicly available datasets and several self-collected datasets, totaling more than 30 sequences. Extensive testing on both X86 and ARM platforms confirms that Super-LIO offers superior efficiency and robustness, while maintaining competitive accuracy. Super-LIO processes each frame approximately 73% faster than SOTA, while consuming less CPU resources. The system is fully open-source and plug-and-play compatible with a wide range of LiDAR sensors and platforms. The implementation is available at: https://github.com/Liansheng-Wang/Super-LIO.git.
Index Terms:
LiDAR-Inertial Odometry, Real-Time Robotics, Resource-Constrained Systems
@@ I Introduction
LiDAR-Inertial Odometry (LIO) is a key component of modern autonomous navigation systems, combining the geometric precision of LiDAR with the high-frequency motion priors from inertial measurement units (IMUs) to achieve outstanding accuracy and robustness. It has been widely deployed in mobile robots [1], unmanned aerial vehicles (UAVs)[2], and autonomous driving[3]. However, deploying LIO on low-power, resource-constrained platforms remains challenging, as it requires balancing accuracy, real-time responsiveness, computational efficiency, and memory footprint, while it ultimately serves as a low-level localization module underpinning higher-level tasks such as exploration [4], language-driven navigation [5], and multi-agent coordination [6], where its value lies in delivering reliable pose estimation with minimal latency and computational cost.
@@
Figure 1: Performance comparison of different LIO methods on X86 and ARM platforms. The left plot shows the average processing time per LiDAR frame, annotated relative to Super-LIO (1x) on each platform, while the right plot shows runtime CPU utilization. Note that the X86 experiments were conducted with 5x playback speed, which results in noticeably higher CPU usage compared to ARM.
Early LiDAR Simultaneous Localization and Mapping (SLAM) systems, such as LOAM [7] and LeGO-LOAM [8], decomposed the SLAM pipeline into feature extraction, frame-to-frame registration, and map maintenance, establishing the foundation for modern LIO. LIO-SAM [9] introduced an incremental pose graph optimization backend based on GTSAM [10], enabling multisensor fusion and global consistency through loop closure. More recent systems, including HAB [11] and Voxel-SLAM [12], incorporated hierarchical global optimization to reduce accumulated drift and improve map quality. While these optimization-based frameworks excel in map fidelity, their iterative solvers scale poorly with map size, resulting in high computational overhead and limited suitability for real-time applications on resource-limited hardware.
In contrast, filtering-based frameworks have emerged as a practical alternative for applications requiring low latency and stable performance. FAST-LIO [13] introduced a tightly coupled LiDAR-IMU fusion architecture using an error-state Kalman filter, while FAST-LIO2 [14] removed explicit feature extraction and leveraged an incremental KD-tree (iKD-tree) for efficient nearest-neighbor queries. Faster-LIO [15] replaced the iKD-tree with a sparse hash-voxel map (iVox) and adopted aggressive parallelization to further improve throughput, albeit at the cost of higher instantaneous computational demand. Point-LIO [16] extended FAST-LIO2 with point-wise state updates, significantly enhancing robustness in high-dynamic scenarios. These systems generally store raw points in the map and use point-to-plane constraints, requiring density control mechanisms to keep K-Nearest Neighbor (KNN) search and planar fitting tractable.
Beyond raw-point maps, VoxelMap [17] introduced a parametric plane-based representation with octree indexing to facilitate local planar feature extraction. iG-LIO [18] drew inspiration from VGICP [19] and stored normalized covariance matrices within hashed voxels, encoding structural information without retaining raw points. While this design reduces storage, its reliance on repeated matrix decompositions and inversions can hinder performance on ARM-based platforms with limited floating-point capability. Liv-GS [20] and GS-LIVO [21] took this further by modeling point-cloud distributions as Gaussian primitives, enabling high-fidelity surface modeling but often requiring GPU acceleration due to their computational complexity.
In this work, we exploit the potential of high-frequency, accurate LiDAR depth measurements and revisit the LIO mapping pipeline to further improve efficiency and free computational resources for downstream tasks. We propose Super-LIO, a robust and efficient system that leverages a sparse hash-voxel structure [15] together with a compact octo-voxel-based representation, termed OctVox. Each voxel is subdivided into eight subvoxels, each storing an incrementally averaged point, enabling global noise suppression and density regularization without increasing hash table complexity. We further introduce a heuristic-guided KNN (HKNN) strategy that leverages precomputed relative neighbor patterns exploiting subvoxel octant symmetry. Evaluations on four public and self-collected datasets (30+ sequences, over 400 minutes) across x86 (AMD 5800H) and ARM (NVIDIA Orin NX) platforms show that Super-LIO matches the accuracy of state-of-the-art systems while significantly reducing computational load (see Fig. 1). The main contributions are:
  • We propose a compact octo-voxel-based map representation (OctVox), where each voxel contains eight incrementally averaged subvoxels, achieving density regularization and global noise suppression without storing raw points.
@@
  • We design a heuristic-guided KNN strategy that exploits subvoxel octant symmetry and unified query templates to improve neighbor quality and query efficiency.
  • We conduct comprehensive evaluations on diverse datasets and hardware platforms, demonstrating that Super-LIO achieves superior runtime efficiency, low CPU usage, and high robustness.
  • We will release the full system to benefit the research community.
II RELATED WORKS
@@
This section reviews prior LIO systems with an emphasis on the design choices in their mapping modules and correspondence search strategies, which are the primary computational bottlenecks in real-time deployment.
LIO frameworks. Optimization-based LIO systems [9, 11, 12, 22] typically aim for global consistency via pose graph optimization or hierarchical bundle adjustment. These methods can achieve low long-term drift but suffer from scalability issues: their iterative nonlinear solvers scale poorly with map size, and runtime grows with the number of accumulated keyframes.
Filtering-based LIO frameworks [14, 15, 16] maintain a bounded state vector and fixed-lag window, ensuring constant per-frame estimation complexity regardless of mapping duration. FAST-LIO2 [14] removed explicit feature extraction and introduced the iKD-Tree for efficient point association. Faster-LIO [15] replaced the iKD-tree with a sparse hash-voxel map, yielding average $\mathcal{O}(1)$ voxel access and improving cache locality by restricting neighbor stencils and scanning contiguous voxel buckets, while throughput is further boosted via parallelization. Point-LIO [16] increased observation bandwidth by updating the state with every valid point-to-plane constraint, improving robustness under aggressive motion. All of these systems follow the error-state Kalman filter formulation on manifolds [23], with extensions such as SI-LIO [24] adopting the invariant EKF [25] to improve linearization consistency. In most cases, the majority of runtime is consumed not by the filter update itself but by the map management and KNN search modules, which is precisely the focus of this work.
Map structures and efficiency. Map structures for LIO differ in their spatial indexing strategies. KD-trees provide $O(\log N)$ query complexity but require frequent rebuilding in incremental mapping, introducing $O(N\log N)$ update cost over time. iKD-Tree [14] mitigates the rebuild overhead by supporting online insertion, but still suffers from $O(\log N)$ amortized update cost and pointer-heavy traversals, which can reduce cache efficiency on modern CPUs. Octrees [26, 27, 22] reduce memory by hierarchical subdivision, but high resolution demands deep trees, increasing pointer traversal cost and reducing cache efficiency.
Hashing-based voxel maps [15] scale well with environment size and support $O(1)$ average voxel access, though KNN efficiency depends on the per-voxel point count. KISS-ICP [28] uses a similar hash-voxel grid but caps the number of stored points per voxel and applies range filtering to bound map size, offering predictable memory and query cost but offering limited control over spatial density distribution. Adaptive-LIO [29] extends this design to multi-resolution hash maps selected dynamically according to scene openness, improving adaptability at the expense of extra memory and neighbor search overhead. Probabilistic voxel methods [19, 18, 30] replace raw points with compact statistical models, avoiding explicit KNN queries but relying on fixed-resolution voxelization, which limits adaptability to varying scene scales. Repeated covariance computations can also suffer from numerical precision loss on ARM-class processors, potentially degrading registration accuracy.
Our approach follows the hashing paradigm but applies an octant subdivision within each voxel, storing one incrementally averaged point per subvoxel to decrease the computation complexity. This ensures constant per-voxel memory usage, enforces point density bounds, and suppresses noise without increasing hash table complexity.
KNN strategies and complexity. In incremental LIO, the cost of KNN search comes from generating candidates and selecting the $k$ nearest ones. KD-trees (including iKD-tree [14]) typically achieve $O(\log N+k)$ query complexity with $O(\log N)$ amortized insertion, but pointer traversal and rebalancing can hurt cache performance. Hash-based voxel maps [15] reduce indexing to $O(1)$ on average, so the runtime mainly depends on the number of points checked within the voxel neighborhood. This number is affected by both the neighborhood size and the per-voxel storage cap; for example, KISS-ICP [28] limits each voxel to $N_{\max}$ points. Faster-LIO [15] restricts the search to only adjacent voxels for speed, but this can reduce neighbor quality in sparse areas.
Our method expands the neighborhood search using a precomputed near-to-far traversal order that exploits octant symmetry within each voxel. With an early-termination rule, this design improves neighbor quality without significantly increasing the number of points that need to be examined.
@@ III-A System Overview
As shown in Fig. 2, Super-LIO is a tightly coupled LiDAR–IMU odometry that follows an iterated error-state Kalman filter (IESKF) formulation [13], while introducing two efficiency-centric modules: (i) OctVox, a sparse hash map that stores at most eight incrementally averaged representatives per voxel for explicit density control and incremental denoising (Sec. III-C); and (ii) a HKNN correspondence module that exploits subvoxel octant symmetry and a near-to-far subvoxel traversal order to restrict neighbor search to a small local neighborhood (Sec. III-D). After IMU-aided de-skewing, each scan is downsampled via a center-based selection strategy that better preserves spatial distribution than standard voxel filters; the selected points are then used both to form scan-to-map point-to-plane residuals for IESKF and to update the OctVox map.
@@ III-B State and Notation
Let $\mathcal{G}$, $\mathcal{L}$, and $\mathcal{I}$ denote the Global, LiDAR, and IMU frames, respectively. The LiDAR–IMU extrinsic transform $\bm{T}_{\mathcal{L}}^{\mathcal{I}}=(\bm{R}_{\mathcal{L}}^{\mathcal{I}},\bm{p}_{\mathcal{L}}^{\mathcal{I}})$ is assumed to be well calibrated and fixed. The system state at time $t$ is defined as Eq.(1).
$\bm{X}_{t}\triangleq\begin{bmatrix}\bm{R}_{\mathcal{I}}^{\mathcal{G}},\ \bm{p}_{\mathcal{I}}^{\mathcal{G}},\ \bm{v}_{\mathcal{I}}^{\mathcal{G}},\ \mathbf{b}_{a},\ \mathbf{b}_{g},\ \bm{g}^{\mathcal{G}}\end{bmatrix}^{\!\top}\in SO(3)\times\mathbb{R}^{15}.$ (1)
We denote estimates with a hat $\hat{(\cdot)}$ and use superscripts $\hat{(\cdot)}^{-}$$\hat{(\cdot)}^{+}$ for priori and posteriori; unless stated otherwise, a bare symbol (e.g., $\bm{X}_{t}$) denotes ground truth.
IMU measurements $\{\bm{a}_{i},\bm{\omega}_{i}\}$ drive propagation at the IMU rate. We use midpoint integration by averaging adjacent samples to reduce discretization error. For each scan anchored at $t_{k}$, raw LiDAR points are first expressed and de-skewed into the IMU frame at the anchor time $t_{k}$:
$\bm{p}^{\mathcal{I}_{k}}_{i}=\bm{R}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}}\!\left(\bm{R}_{\mathcal{L}}^{\mathcal{I}}\bm{p}^{\mathcal{L}}_{i}+\bm{p}_{\mathcal{L}}^{\mathcal{I}}\right)+\bm{p}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}},\ t_{i}\!\in\![t_{k-1},t_{k}],$ (2)
where $(\bm{R}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}},\,\bm{p}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}})$ denotes the relative IMU motion from $t_{i}$ to $t_{k}$ obtained by propagation and subsequent interpolation between IMU samples, for example using a constant–velocity model [7], a constant–acceleration model [14], or a higher-order scheme [31]. This convention is used throughout: de-skewed points are retained in the IMU frame $\mathcal{I}_{k}$ and mapped to the world using Eq.(3) when needed for association or map update.
$\bm{p}^{\mathcal{G}}_{i}=\bm{R}_{\mathcal{I}}^{\mathcal{G}}(t_{k})\,\bm{p}^{\mathcal{I}_{k}}_{i}+\bm{p}_{\mathcal{I}}^{\mathcal{G}}(t_{k}).$ (3)
Selected points then form scan-to-map point-to-plane residuals for the IESKF observation update [13], while the compact OctVox map supplies correspondence targets.

### v2

@@ Abstract
LiDAR-Inertial Odometry (LIO) is a foundational technique for autonomous systems, yet its deployment on resource-constrained platforms remains challenging due to computational and memory limitations. We propose Super-LIO, a robust LIO system that demands both high performance and accuracy, ideal for applications such as aerial robots and mobile autonomous systems. At the core of Super-LIO is a compact octo-voxel-based map structure, termed OctVox, that limits each voxel to eight subvoxel representatives, enabling strict point density control and incremental denoising during map updates. This design enables a simple yet efficient and accurate map structure, which can be easily integrated into existing LIO frameworks. Additionally, Super-LIO designs a heuristic-guided KNN strategy (HKNN) that accelerates the correspondence search by leveraging spatial locality, further reducing runtime overhead. We evaluated the proposed system using four publicly available datasets and several self-collected datasets, totaling more than 30 sequences. Extensive testing on both X86 and ARM platforms confirms that Super-LIO offers superior efficiency and robustness, while maintaining competitive accuracy. Super-LIO processes each frame approximately 73% faster than SOTA, while consuming less CPU resources. The system is fully open-source and compatible with a wide range of LiDAR sensors and computing platforms. The implementation is available at: https://github.com/Liansheng-Wang/Super-LIO.git.
@@ I Introduction
Figure 1: Performance comparison of different LIO methods on X86 and ARM platforms. The left plot shows the average processing time per LiDAR frame, annotated relative to Super-LIO (1x) on each platform, while the right plot shows runtime CPU utilization. Note that the X86 experiments were conducted with 5x playback speed, which leads to higher CPU usage compared to ARM.
LiDAR-Inertial Odometry (LIO) is a core module in autonomous navigation systems, fusing LiDAR geometry with inertial motion priors for accurate and robust pose estimation. It has been widely deployed in mobile robots, UAVs, and autonomous driving [AR_ugv, super, AotoDrive]. As robotic tasks grow in complexity, LIO increasingly operates as a foundational layer for higher-level autonomy, such as exploration [epic], language-driven navigation [openfly], and multi-agent coordination [uav_swarm], where limited onboard computational resources must be shared among multiple modules. This competition for resources, especially on embedded or low-power platforms, makes it challenging for LIO to maintain real-time performance and high accuracy.
Early LiDAR-based Simultaneous Localization and Mapping (SLAM) systems, such as LOAM [LOAM] and LeGO-LOAM [lego], decomposed the pipeline into feature extraction, scan-to-map registration, and map maintenance, establishing the foundation for LIO frameworks. LIO-SAM [lio_sam] further incorporates a pose-graph optimization backend to achieve loop closure and maintain global consistency. Later systems such as HBA [hba] and Voxel-SLAM [Voxel-slam] employ hierarchical optimization combined with bundle adjustment to perform global map refinement after the front-end odometry stage, further improving mapping quality. While these approaches achieve higher map fidelity, their optimization backends introduce additional computational stages, leading to higher per-frame latency and reducing their suitability for latency-sensitive deployment on resource-constrained platforms [fast_lio2, PointLIO, faster_lio].
Filtering-based LIOs offer lightweight estimation and minimal latency, where the dominant computational cost lies in point-cloud registration due to extensive nearest-neighbor search. This cost and its consistency are largely determined by the underlying map structure, which governs both the number and spatial locality of candidate neighbors. Incrementally updatable maps such as the iKD-Tree [fast_lio2] and hash-voxel grids like iVox [faster_lio] accelerate point insertion and neighbor queries, but their runtime can still vary with factors such as local point density or voxel discretization. KISS-ICP [Kiss-icp] adopts a similar hash-voxel design that limits stored points per voxel to control memory usage. Although such structures improve efficiency, they rely on voxel-level thresholds rather than explicitly enforcing spatially uniform point distribution.
To overcome the density-related variability of existing voxel maps, we introduce OctVox, a compact hash-based octo-voxel representation that explicitly regularizes spatial density. Each voxel is subdivided into eight subvoxels, each maintaining an incrementally averaged representative point. This design provides bounded storage, uniform coverage, and a compact geometric representation, allowing a smaller yet well-structured point set to support accurate odometry while reducing computational and memory costs.
While OctVox stabilizes the map structure, achieving high-quality correspondences still requires searching beyond the small fixed neighborhoods used in prior voxel-based methods. Voxel-based neighbor queries [faster_lio, ig_lio] incur rapidly increasing cost as the search range expands. Building upon the ordered subvoxel layout of OctVox, we develop a heuristic-guided K-nearest-neighbor strategy (HKNN) that efficiently searches larger candidate regions with early termination, improving neighbor quality without incurring excessive overhead.
We integrate OctVox and HKNN into a unified filtering-based LIO framework, Super-LIO. Extensive experiments on four public datasets and multiple self-collected sequences across x86 and ARM platforms show that Super-LIO achieves accuracy comparable to state-of-the-art methods while reducing runtime and CPU usage (see Fig. 1). The main contributions are summarized as follows:
  • We propose a compact hash-based octo-voxel map representation (OctVox) that explicitly regularizes spatial density, achieving uniform coverage and noise suppression.
@@
  • We design a heuristic-guided KNN strategy (HKNN) that leverages the structured subvoxel layout of OctVox to improve neighbor quality and search efficiency.
  • We integrate OctVox and HKNN into a unified filtering-based LIO system (Super-LIO) and validate it on diverse public and self-collected datasets across both x86 and ARM platforms, showing accuracy comparable to state-of-the-art methods with lower runtime and CPU usage.
  • We release the complete system to benefit the research community.
II Related Work
@@
This section reviews prior LIO systems with a focus on map structures and correspondence search strategies, which dominate the computational cost in real-time deployment.
Map structures and efficiency. Early LIO systems such as LIO-SAM [lio_sam] and FAST-LIO [fast_lio] organize local maps using KD-Trees for correspondence search, whose maintenance cost becomes non-negligible due to frequent rebuilding in long-term incremental mapping. FAST-LIO2 [fast_lio2] addresses this limitation by introducing the iKD-Tree, enabling efficient online updates without full reconstruction. Building upon the iKD-Tree framework, Point-LIO [PointLIO] further adopts point-wise state updates to increase update bandwidth and improve robustness under high-dynamic motion.
Faster-LIO [faster_lio] replaces tree-based indexing with a sparse hash voxel map, reducing insertion overhead and yielding more predictable access patterns. However, the number of points stored per voxel is not explicitly controlled, causing correspondence search cost to depend on local point density and potentially increase in dense regions. KISS-ICP [Kiss-icp] adopts a similar hash-voxel structure but caps the number of points per voxel to bound indexing cost, while still lacking explicit spatial density regulation. Adaptive-LIO [adaptive-lio] adjusts voxel resolution according to scene openness, at the expense of increased structural complexity and higher neighborhood query cost.
Voxel-based statistical methods offer an alternative representation by replacing raw points with compact statistical models VGICP [vgicp] estimates voxel-wise means and covariances to replace explicit nearest-neighbor queries with probabilistic distance evaluations. iG-LIO [ig_lio] builds upon this idea [vgicp] and integrates it into a tightly coupled LIO framework, improving overall efficiency. However, its statistical modeling requires environment-specific parameter tuning, and its computational demand is relatively high on embedded processors. GLIM [GLIM] adopts a similar formulation but relies on GPU acceleration, which requires additional hardware support.
Octrees [ufomap, slict, i-octree] reduce memory through hierarchical subdivision, but high-resolution settings lead to deeper tree structures and longer traversal paths during neighbor lookup in high-rate LIO. MGM-LIO [MGM-LIO] employs multi-scale Gaussian maps, but its mapping and update procedures remain computationally heavy for real-time use on mobile robots. Recent Gaussian-splatting-based LiDAR mapping methods [Liv-GS, GS-LIVO] rely on GPU acceleration, making them less suited for lightweight platforms.
Across these designs, spatial density is generally unmanaged, or addressed indirectly through statistical modeling. To address this, we introduce OctVox, which enforces subvoxel-level density regularization and provides a compact, structured representation for efficient mapping.
KNN strategies and complexity. In scan-to-map registration, correspondence estimation relies on repeated nearest-neighbor queries and constitutes a major runtime bottleneck.
KD-tree-based methods [fast_lio2] perform KNN search through recursive space partitioning, whose backtracking behavior makes the search cost sensitive to point distribution and local geometry.
Hash-voxel-based approaches such as iVox [faster_lio] achieve constant-time voxel access via hashing, but the KNN cost depends on the number and spatial distribution of points within the queried voxels. To bound this cost, searches are typically restricted to fixed neighboring voxel stencils; however, due to spatial discretization, true nearest neighbors may fall outside the stencil, leading to a resolution- and density-dependent trade-off between efficiency and correspondence quality. Methods such as iG-LIO [ig_lio] and MGM-LIO [MGM-LIO] follow similar designs with per-voxel statistical models, but still rely on fixed voxel neighborhoods and incur additional probabilistic evaluation cost.
Building on the structured subvoxel layout of OctVox, we develop an HKNN strategy that leverages spatial priors to expand the effective search region while avoiding unnecessary point evaluations, enabling higher-quality correspondences at lower computational cost.
@@ III-A System Overview
As shown in Fig. 2, Super-LIO is a tightly coupled LiDAR-Inertial Odometry that follows an iterated error-state Kalman filter (IESKF) formulation [fast_lio], while introducing two efficiency-centric modules: (i) OctVox, a sparse hash map that stores at most eight incrementally averaged representatives per voxel for explicit density control and incremental denoising (Sec. III-C); and (ii) HKNN, a correspondence module that uses precomputed traversal lists and subvoxel symmetry to accelerate neighbor search within a small local neighborhood (Sec. III-D). After IMU-aided de-skewing, each scan is downsampled via a center-based selection strategy that better preserves original measurements than standard voxel filters; the selected points are then used both to form scan-to-map point-to-plane residuals for IESKF and to update the OctVox map.
@@ III-B State and Notation
Let $\mathcal{G}$, $\mathcal{L}$, and $\mathcal{I}$ denote the Global, LiDAR, and IMU frames, respectively. The LiDAR-IMU extrinsic transform $\mathbf{T}_{\mathcal{L}}^{\mathcal{I}}=(\mathbf{R}_{\mathcal{L}}^{\mathcal{I}},\mathbf{p}_{\mathcal{L}}^{\mathcal{I}})$ is assumed to be well calibrated and fixed. The system state at time $t$ is defined as Eq.(1).
$\mathbf{X}_{t}\triangleq\begin{bmatrix}\mathbf{R}_{\mathcal{I}}^{\mathcal{G}},\ \mathbf{p}_{\mathcal{I}}^{\mathcal{G}},\ \mathbf{v}_{\mathcal{I}}^{\mathcal{G}},\ \mathbf{b}_{a},\ \mathbf{b}_{g},\ \mathbf{g}^{\mathcal{G}}\end{bmatrix}^{\!\top}\in SO(3)\times\mathbb{R}^{15}.$ (1)
We denote estimates with a hat $\hat{(\cdot)}$ and use superscripts $\hat{(\cdot)}^{-}$$\hat{(\cdot)}^{+}$ for priori and posteriori; unless stated otherwise, a bare symbol (e.g., $\mathbf{X}_{t}$) denotes ground truth.
IMU measurements $\{\mathbf{a}_{i},\bm{\omega}_{i}\}$ drive propagation at the IMU rate. We use midpoint integration by averaging adjacent samples to reduce discretization error. For each scan anchored at $t_{k}$, raw LiDAR points are first expressed and de-skewed into the IMU frame at the anchor time $t_{k}$:
$\bm{p}^{\mathcal{I}_{k}}_{i}=\mathbf{R}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}}\!\left(\mathbf{R}_{\mathcal{L}}^{\mathcal{I}}\bm{p}^{\mathcal{L}}_{i}+\mathbf{p}_{\mathcal{L}}^{\mathcal{I}}\right)+\mathbf{p}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}},\ t_{i}\!\in\![t_{k-1},t_{k}],$ (2)
where $(\mathbf{R}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}},\,\mathbf{p}_{\mathcal{I}_{i}}^{\mathcal{I}_{k}})$ denotes the relative IMU motion from $t_{i}$ to $t_{k}$ obtained by propagation and subsequent interpolation between IMU samples, for example using a constant-velocity model [LOAM], a constant-acceleration model [fast_lio2], or a higher-order scheme [dlio]. De-skewed LiDAR points are expressed in the IMU frame $\mathcal{I}_{k}$ and transformed to the global frame via Eq. (3) when needed for association or map updates.
$\bm{p}^{\mathcal{G}}_{i}=\mathbf{R}_{\mathcal{I}}^{\mathcal{G}}(t_{k})\,\bm{p}^{\mathcal{I}_{k}}_{i}+\mathbf{p}_{\mathcal{I}}^{\mathcal{G}}(t_{k}).$ (3)